Grade 12th passMechanics

a car is moving at a constant speed of 40 km per hour along a straight road which headds towards a large vertical wall and makes a sharp 90 degree turn by the side of a wall a fly flying at a constant speed of 100 km per hour starts from the wall towards the car at an instant when the car is no 20 kilometre away flies until it with the glass plane of the car and returned to the wall at an SMS Fitbit contribute to this line between the car and wolf till the car makes the 90 degree turn what is the total distance the fly has travelled during this period how many teeth has it made between the car and the wall

Sachin
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago

distance travelled = velocity x time

 = 100 x t

Here, t is the time that the car takes to complete 20 km with 40 km/h = 1/2 hour

=> distance = 100/2 = 50 km

b)   No. of trips = infinite , as distance between car and wall tends to 0, the no. of trips of fly tends to infinity.